Calculate the pressure of an object having a force of 42lb with an area of 6.0 inches2


1
Expert's answer
2021-03-14T19:13:48-0400

By the definition of the pressure, we have:


"P=\dfrac{F}{A}=\dfrac{42\ lb}{6.0\ in^2}=7\ \dfrac{lb}{in^2}."